### v2

# Sensitive Content Classification in Social Media: A Holistic Resource and Evaluation

Abstract
The detection of sensitive content in large datasets is crucial for ensuring that shared and analysed data is free from harmful material. However, current moderation tools, such as external APIs, suffer from limitations in customisation, accuracy across diverse sensitive categories, and privacy concerns. Additionally, existing datasets and open-source models focus predominantly on toxic language, leaving gaps in detecting other sensitive categories such as substance abuse or self-harm. In this paper, we put forward a unified dataset tailored for social media content moderation across six sensitive categories: conflictual language, profanity, sexually explicit material, drug-related content, self-harm, and spam. By collecting and annotating data with consistent retrieval strategies and guidelines, we address the shortcomings of previous focalised research. Our analysis demonstrates that fine-tuning large language models (LLMs) on this novel dataset yields significant improvements in detection performance compared to open off-the-shelf models such as LLaMA, and even proprietary OpenAI models, which underperform by 10-15% overall. This limitation is even more pronounced on popular moderation APIs, which cannot be easily tailored to specific sensitive content categories, among others. Equal contribution.
[…]
Sensitive Content Classification in Social Media:
A Holistic Resource and Evaluation
Dimosthenis Antypas*1, Indira Sen*2, Carla Perez-Almendros1 Jose Camacho-Collados1, Francesco Barbieri 1 Cardiff NLP, Cardiff University, United Kingdom 2 University of Mannheim 1 {AntypasD,CamachoColladosJ,perezalmendrosc}@cardiff.ac.uk 2 indira.sen@uni-mannheim.de
Disclaimer
Due to the nature of the subject studied in this work, this paper contains sensitive and potentially offensive language. Reader discretion is advised.
1 Introduction
Consider the case of a researcher or a data analyst who needs to filter sensitive content from a large dataset. Such a task is crucial to ensure that data shared or analysed does not include harmful or inappropriate material. One might initially consider using external tools like Perspective [^1] https://perspectiveapi.com/ or OpenAI moderation APIs [^2] https://platform.openai.com/docs/guides/moderation to assess and filter sensitive content. However, this approach often falls short, presenting important limitations for an effective identification of inappropriate content online Udupa et al. (2023). For instance, they usually offer limited customisation capabilities (e.g., how can the model be improved if it fails on specific domains or keywords?), and limited sensitive categories coverage (lacking in detecting self-harm Uban and Rosso (2020) for example). Finally, these tools rely on external servers, which raises concerns about data privacy and security Oseni et al. (2021); Gupta et al. (2023).
Alternatively, one might consider using existing datasets and open source models for sensitive content detection. This could be a viable option if the primary focus was on detecting toxic language, given the abundance of resources available in this area. However, if the goal extends to identifying additional sensitive categories such as sexually explicit content, drugs, self-harm, and spam, the situation becomes more challenging. Data on these less-explored categories is limited and sometime outdated. For instance, those categories could be covered on dataset that are biased Wiegand et al. (2019), old or inaccessible even in an anonymized manner Tadesse et al. (2019); Sawhney et al. (2018), too small-scaled or rely on an handful of keywords to extract the data Ding et al. (2016). This limited approach can result in incomplete or less accurate detection of sensitive content.
Existing solutions either require sending data to external servers or fail to address the full spectrum of sensitive content categories. In response to these challenges, this paper proposes a new holistic approach: a unified dataset for detecting sensitive text across a broad range of categories, including (1) conflictual language, (2) profanity, (3) sexually explicit material, (4) drug-related content, (5) self-harm, and (6) spam [^3] We use social media platform user guidelines to be sure to have reliable guidelines.. This dataset can then be used for both evaluation and fine-tuned model to all these categories under a single framework.
Our approach involves collecting and re-annotating data to ensure consistent quality across sensitive classes. The alternative of putting together a collection of existing datasets to create one single dataset would include several limitations as (1) there would be a bias towards the data distribution (different retrieval strategies, topics, source platforms), (2) annotation guidelines and quality would differ, (3) each text would include only one sensitive dimension (even if the text includes multiple sensitive categories).
In short, we propose a holistic approach when it comes to sensitive content moderation in social media, overcoming common shortcomings of previous works and providing the following contributions:
  • We introduce the X-Sensitive dataset, manually annotated and tailored for social media content, featuring multiple categories and designed to be resilient against keyword and domain shifts.
  • Study the interplay between sensitive categories and how these categories vary across different annotator demographics.
  • Analyse the performance of current models: The best results are achieved from large language models (LLMs) with 8 billion parameters, fine-tuned on our dataset. However, smaller language models [^4] Pre-trained on social media language. (355 million parameters) show only about 2% less accuracy.
  • Comparison with off-the-shelf LLMs: We find that readily available LLMs, such as gpt-4o, under-perform by 10-15% compared to fine-tuned models, highlighting the value of bespoke training on specialised datasets.
The X-Sensitive dataset, as well as the best performing models built upon it, are made openly available. X-Sensitive is available at https://huggingface.co/datasets/cardiffnlp/x_sensitive. Best multi-label and binary models are available at https://huggingface.co/cardiffnlp/twitter-roberta-large-sensitive-multilabel and https://huggingface.co/cardiffnlp/twitter-roberta-large-sensitive-binary, respectively.
2 Related Work
Our current work aims to bridge the gap between current academic research in content moderation and the needs of content moderators in realistic scenarios. While hate speech and toxic language are widely studied in NLP, there is little research on other types of sensitive content that platforms seek to detect and moderate, such as sexually explicit content or content about illicit substances Arora et al. (2023). To that end, our work in situated at the intersection of NLP research on harmful language detection and research on platform governance and content moderation.
2.1 Automatic Detection of Harmful Language
Hate speech Detection.
Automatic detection of hate speech, and related social constructs like offensive and toxic language, is an active area of research in NLP Fortuna and Nunes (2018); Poletto et al. (2021). However, there are several challenges, not least the lack of high quality datasets for studying such phenomena Vidgen and Derczynski (2020).
Self-harm and Suicidal Content Detection.
Chancellor et al. (2016b) identify communities with self-harm related content, while Tejaswini et al. (2024) also look into related behaviors such as depression. Previous research has also looked into suicidal content detection Coppersmith et al. (2018) and general self-harm Un Nisa and Muhammad (2021). There are generally several ethical challenges associated with studying mental health conditions, including self-harm and suicidal ideation Chancellor et al. (2019).
Illicit Substance Abuse.
Past research has looked into automated approaches for discussions of illegal or banned substances, including drugs Buntain and Golbeck (2015); Lavanya and Sasikala (2022); Simpson et al. (2018).
Sexually Explicit Content.
Research has also focused on developing automated systems to detect sexually explicit content Barrientos et al. (2020), address sexual harassment Chowdhury et al. (2019), and identify sexualised cyberbullying Basu et al. (2021).
Spam Detection.
Automatic Spam detection is widely studied in NLP as well as computer security communities. Typical automation techniques rely on expert-annotated training data used to train machine learning models Hussain et al. (2019). However, like the other categories spam detection has rarely been studied in the context of other types of problematic content, with Founta et al. (2018) being an exception.
2.2 Content Moderation and Platform Governance
Platforms on the internet, such as web and social media sites, often employ mechanisms to curate their content and reduce problematic or harmful content through content moderation (CM). CM can take many forms, from commercial content moderation outsources to underpaid moderators in the Global South Roberts (2019) to artisanal solutions, some of which are led by volunteers Caplan (2018). Yet as content grows, platforms turn towards automated methods, often Artificial Intelligence (AI) based techniques either solve or ameliorate their moderation problem Gorwa et al. (2020).
However, the question remains on how much of this detection is automatable? Gillespie (2020). There are not only several technological limitations (e.g. the dearth of AI methods for non-English content Vidgen and Derczynski (2020)) but also political challenges (e.g., who gets to decide what is harmful? Fleisig et al. (2024)) and challenges at the nexus of technology and politics (e.g., how do we aggregate the potentially divergent judgements of whether something is harmful? Fan and Zhang (2020); Gordon et al. (2022)). On the other hand, platform studies researchers have studied which types of technological solutions, including AI-based tools, would facilitate the work of content moderators while also establishing some of the tensions of the whole practice of content moderation. However, it is unclear if those proposing technological solutions for CM are basing their solutions on the requirements of content moderators.
Categories of Sensitive Content. Several researchers have attempted to categorise what counts as ‘sensitive’ content on web and social media platforms, i.e., content that requires moderation Jiang et al. (2020); Scheuerman et al. (2021).
We address one of the many challenges of automatic content moderation — lack of benchmark datasets for measuring understudied categories of problematic content like discussion related to self-harm and illicit substances, particularly drugs. We also provide a holistic benchmark of both these aforementioned understudied categories as well as widely studied categories like profanity, allowing researchers to model the associations between different types of sensitive content.
3 X-Sensitive Dataset
In order to study sensitive content in X, we construct a new dataset, X-Sensitive. As a first step, we conceptualise a topic taxonomy based on community guidelines from several social media platforms.
3.1 Taxonomy
We use the community guidelines of various social media platforms to ground our taxonomy Scheuerman et al. (2021). Using iterative coding, we refine, merge, and fix 5 broad categories and 7 specific subcategories of sensitive content which are mapped to rules in community guidelines. Our final categories and their definitions are:
Drugs.
Content that encourages, promotes or glorifies the use of regulated drugs. Also applicable to content that mentions sales, purchases, or the act of obtaining or trying to obtain regulated drugs.
Sexually Explicit Content (Sex).
Pornographic or other types of sexual content. We collect and download 50+ textual abusive language datasets from hatespeechdata.com. We then use the Perspective API to label these datasets with the ‘sexually explicit’ endpoint and then use the labeled data for finetuning the XLM-T sexually explicit content classifier.
Hate speech.
Attacks against protected attributes like race, color, caste, ethnicity, national origin, religion, sexual orientation, gender identity, disability, or veteran status, immigration status, socio-economic status, age, weight or pregnancy status.
Other conflictual language.
Attacks based on other categories or without any mention of the categories mentioned.
Profanity.
Language containing slurs and profanity even if they are not directed towards a specific entity.
Self-harm.
Posts depicting, promoting or glorifying violence or harm against oneself, such as eating disorders or suicide.
Spam.
Irrelevant content that is unsolicited; or content that aims to drive traffic or attention from a conversation on the platform to entities outside the platform.
3.2 Tweet Collection
Typically previous work on sensitive content detection, particularly hate speech detection, uses a small set of keywords to collect data, which may lead to limited coverage of the resultant datasets Ousidhoum et al. (2021). To tackle this problem, we utilise a keyword expansion technique combining word embeddings Mikolov et al. (2013), trained on tweets Pennington et al. (2014), for keyword expansion and clustering for controlling the expanded sets. The specific algorithm is described as follows:
Algorithm 1 Keyword List Expansion Technique using Word Embeddings
1:Input: Seed list $\{w_{1},w_{2},\ldots\}$
2:Output: Expanded keyword list
3:Start with a seed list $\{w_{1},w_{2},\ldots\}$
4:Cluster keywords’ vectors into $k$ clusters
5:Check and select the relevant clusters
6:Compute the dot product $\mathbf{v}_{i}\cdot\boldsymbol{\mu}_{c}$ for each word vector $\mathbf{v}_{i}$ and each cluster mean vector $\boldsymbol{\mu}_{c}$
7:Find $n_{1}$ words that are closest to the mean of each cluster
8:for each new word do
9:     Find the closest $n_{2}$ words
10:end for
Due to the variety of the categories in our sensitive content category, we look at multiple sources for seed lists (Appendix A, Table 3). The conceptual similarities between profanity, sexually explicit content, and conflictual content as well as the existence of lists that collect keywords related to these three types of discourse, make us opt for a unified seed word list covering these three categories.
For self-harm and spam, unigrams are not as informative as they are for other categories. Therefore for the former, we look at past research on eating disorders and suicidal ideation Chancellor et al. (2016b, 2021) and obtain phrases (ngrams with a high TF-IDF score) from the Reddit data used to train the self-harm phase 1 classifier. For spam, we use the dataset from Founta et al. (2018), particularly the tweets that were labelled containing spam and obtain ngrams from there using a similar method. We manually assess each of the keywords for all categories and remove low precision words like ‘snow’ for drugs. While snow may refer to cocaine in some contexts, most tweets containing it do not use it in that sense. After this manual inspection, we apply our cluster-based keyword expansion technique. We again manually assess the keywords and include only those that are relevant to the category. The final statistics of our keywords are listed in Appendix A, Table 4.
3.3 Annotation
Each entry of the dataset was annotated by at least three coders, where each coder had to answer with yes, no, or not sure if the tweet contained any of the sensitive classes. Specifically, for the case of conflictual language the annotators were asked to select whether the tweet contained hate speech or any other form of conflictual language. This approach aimed for a more fine grained classification of conflictual language. However, due to low agreement between coders we opted to merge the categories "Hate Speech" and "Other Conflictual Language" into a single class Conflictual Language.
A label was assigned to a tweet if at least one annotator answered yes and there was no direct opposition from the rest of the coders (i.e. the rest of the coders answered yes or not sure). We refrained from utilising a majority rule in order to create a more realistic and challenging dataset while also weighting the recall of potentially sensitive content higher.
The coders who worked on this task were selected and filtered through the Prolific.co platform based on their fluency in English. The annotation was performed through an interface created with qualtricsXM [^5] The annotation guidelines can be found in Figure 3, Appendix B.. The coders were also provided with 15 examples of already annotated tweets to help them better understand the task. Finally, we utilised several filters to ensure a high quality of annotations. First, we included a set of test questions randomly inserted in the task which were used to filter out low quality coders. Additionally, coders that finished the task too quickly or provided low quality answers (for example always selecting the same answer) were excluded.
Overall 523 coders from various demographic backgrounds took part in the annotation process. We assessed the quality of the annotation by utilising Krippendorff’s Alpha (Alpha) Krippendorff (2011). The annotators achieve 0.49 Alpha when considering all available classes and 0.56 Alpha when considering only the presence of sensitive content or not. The scores are in line or better with previous similar studies on toxic and sensitive content Muralikumar et al. (2023); Lima et al. (2024).
It is interesting to note that when looking at subgroups of annotators based on their demographics we observe higher agreement between specific groups, mainly younger (0.51 Alpha in multi-label setting for people 39 old and younger) and non-binary people (0.82 Alpha in the binary setting). More detailed results can be found in Appendix B, Tables 6 and 5).
Looking in more detail on how different demographics annotate examples a trend is noticed where younger coders and non-binary annotators tend to be more sensitive to the content and are more likely to flag a tweet as sensitive (Appendix B, Tables 7 and 8).
The discrepancies in agreement between groups indicate the inherent difficulty of the task while also providing evidence of a greater coverage of sensitive content within X-Sensitive.
3.4 Statistics
X-Sensitive contains a total of 8,000 tweets all related to sensitive content with 49% of them labelled as one or more of the six sensitive classes available making it a challenging dataset. On average tweets flagged as sensitive are assigned 1.4 labels with maximum assigned labels to a single tweet being 4.
Figure 1: Overlap of classes.
Our dataset displays a skewed distribution of classes as seen if Table 1 with profanity being the most populated class present (30.4%). This uneven distribution represents a realistic representation of sensitive content in social media as seen in previous similar studies Beknazar-Yuzbashev et al. (2022) where is estimated a 5% - 7% of content displayed is inappropriate, making X-Sensitive ideal for usage in real world applications.
At the same time profanity being the most frequent class is also expected. Due to the multi-label nature of the dataset, we expect high overlap between profanity and other classes as seen in Figure 1. Particularly there is a high overlap between profanity tweets and those labelled as sexual explicit content, and conflictual.
In general differences between the classes are revealed even when looking at basic statistics such as the average length of tweets and the presence of emojis in them. As seen in Table 1 tweets labelled as spam tend to be longer on average and include a higher number of emojis, characteristics frequently found on spam messages Robinson and Mago (2022). Similarly, a higher usage of emojis is observed in tweets flagged as sexual explicit content, as specific emojis are often used as representation of sexual acts Thomson et al. (2018). Furthermore, when examining the top terms of each class based on lexical specificty scores Camacho-Collados et al. (2016), a clear distinction between the classes is observed, which also serves as a sanity check for the quality of our dataset.
4 Experimental Setting
In this section, we set out the common experimental framework which serve for the basis of the evaluation.
4.1 Data and Settings
To evaluate X-Sensitive and establish baselines of it’s difficulty we establish two distinct settings: binary and multi-label classification. In the binary setting, tweets will be classified into one of two categories, distinguishing between sensitive and not sensitive content. This approach simplifies the classification process, focusing on the presence or absence of sensitive characteristics in general. In the multi-label setting, tweets can belong to multiple sensitive categories simultaneously, allowing for a more finegrained analysis that captures the complexity of the content. This dual approach enables a comprehensive evaluation of our dataset’s versatility and the classifier’s robustness in handling varying degrees of complexity in sensitive content detection.
Category L Emo % Top Terms
Conflictual 188.67 0.26 17.3
fucking racist
nigga white shut
Profanity 173.05 0.43 30.4
fucking shit
bitch fuck as
Sex 160.79 0.66 9.7
cock pussy
dick horny cum
Drugs 155.62 0.21 3.9
drug weed cbd
thc mushroom
Self-harm 166.35 0.44 3.0
suicide suicidal
attempt commit
ideation
Spam 200.30 1.12 3.4
dm project airdrop
solana solanaairdrop
Not Sensitive 176.39 0.32 51.2
physically depressing
triggering mental
depression
Overall 174.77 0.37
Table 1: General lexical statistics for each class. The averages of the length of tweet, emojis count are reported. The distribution of each class along with the top five terms based on their lexical specificity are also displayed.
For both settings we use a split the dataset in train/validation/test sets of 6,000/1,000/2,000 tweets while ensuring that the distribution of classes is similar in each split. To investigate the generalisability capabilities of the models, an additional constraint check is enforced where we ensure that approximately half of the test set, 1,016 tweets, do not share any of the keywords used for collection with tweets from the train set.
4.2 Comparison Systems
For the evaluation we are interested in comparing three types of approaches: finetuning on the same dataset (Section 4.2.1), LLMs with in-context learning either zero- or few-shot (Section 4.2.2), and out of the box content moderation systems (Section 4.2.3). All the systems are clearly not fully comparable, but our dataset can serve as the basis for establishing this basic ground comparison.
4.2.1 Finetuning
We evaluate three distinct models tailored for various applications, including general-purpose and those specialized for social media, each differing in size for our fine-tuning experiments. The large version of RoBERTa Liu (2019) is tested in order to assess the performance of smaller, non specialised, masked language models on our dataset. Time-lm, tlm, Loureiro et al. (2022), a RoBERta based model trained on a large X corpus of 154 million tweets is also evaluated to assess the performance of specialised models on social media. The two models are fine-tuned using the implementations provided by Hugging Face Wolf et al. (2020) and optimising hyper parameters (learning rate, training epochs, warmup steps) is conducted using Ray Tune Liaw et al. (2018) [^6] Details of the models used can be found in Appendix C.. Finally, the 8 billion version of Llama-3, Llama3-8b, AI@Meta (2024) is also fine-tuned on our dataset by utilising quantisation and PEFT Liu et al. (2021); Mangrulkar et al. (2022) explore the capabilities of more recent and larger-scale models.
4.2.2 Zero- and Few-shot
In order to assess the zero/few-shot capabilities of large language models in our dataset, we compare four models of different sizes and architectures.
Llama3: The 8 and 70 billion instruct versions of Llama3 are tested. These models are designed to follow user instructions more effectively, allowing us to assess how well they adapt in settings where training data is limited or not available.
chat-gpt-3.5-turbo (chat-gpt): from OpenaAI,  [^7] https://openai.com/chatgpt an encoder/decoder model with approximately 175 billion parameters Brown et al. (2020).
gpt-4o: the currently latest model from OpenAI which significantly outperforms it’s predecessor.
Assessing the performance in zero- and few-shot settings, help us to exlore the capabilities and limitations of these large language models for sensitive content detection.
Training Model Binary multi-label Conflictual Profaninty Sex Drugs Selfharm Spam Not sens.
fine-tuned RoBERTa 82.4 64.7 60.6 88.9 81.6 52.3 34.3 52.0 83.3
tlm 84.4 67.7 59.6 88.8 84.3 48.9 50.6 59.1 82.4
llama3-8b 85.6 69.8 61.7 90.6 85.8 53.9 50.6 61.2 85.1
Zeroshot llama3-8b 75.0 52.2 53.5 69.8 70.0 39.2 35.5 21.6 75.6
llama3-70b 76.5 57.4 54.5 79.4 74.3 55.0 32.8 42.1 63.5
chat-gpt 60.0 63.2 49.0 60.0 71.0 57.0 41.0 37.0 69.0
gpt-4o 75.7 64.9 62.2 82.9 84.0 64.9 53.2 26.2 81.1
Fewshot llama3-8b 74.9 53.2 43.9 73.3 74.8 49.0 18.5 43.0 70.1
llama3-70b 79.2 63.0 62.2 82.8 78.9 61.5 32.8 53.9 69.1
chat-gpt 71.0 64.0 59.0 84.0 83.0 52.0 27.0 48.0 72.0
gpt-4o 83.3 67.9 63.4 85.7 81.7 61.1 41.9 64.8 76.9
Out of the box Systems llama-g 55.0 - 16.1 - 75.5 - 43.6 - 68.2
openai-m 72.0 - 63.1 - 73.0 - 46.3 - 75.9
Perspective 70.0 - 64.0 89.0 81.0 - - 53.0 44.0
Table 2: Macro F1 scores for fine-tuned and zero-/few-shot models are reported in both binary and multi-label settings. Additionally, the F1 scores for each class in the multi-label setting are provided. For out-of-the-box systems, we report the F1 scores in the binary setting and, when available, the F1 scores achieved in each class.
4.2.3 Out of the box Systems
The need for detecting sensitive or harmful content has led to several companies to develop their own models, which are made publicly available. In order to highlight the relevance of existing models for this task, we selected three popular specialised systems.
Google’s Perspective API (Perspective) Google (2023) is a tool developed to detect and score various attributes of text, such as toxicity, and acts as a baseline performance of a production-ready API. In total Perspective provides scores for 16 different categories but in our use case we focus only on 12 of them that fit our taxonomy better [^8] Detailed taxonomy can be found in Appendix C.
OpenAI’s moderation API (openai-m) is an endpoint tailored for content moderation. It classifies content into 18 potentially sensitive categories, 15 of which we map to our own taxonomy.
Meta-Llama-Guard-2-8B (llama-g) Inan et al. (2023) is a specialised version of LLama-3 that aims to classify content based on a safety risk taxonomy of 11 harm categories Vidgen et al. (2024)\@footnotemark. For our use case we consider only 5 of the categories that correspond better to the taxonomy used in X-Sensitive. Specifically we consider: "Hate" for Conflictual, "Suicide & Self-Harm" for self harm; and "Sexual Content", "Sex-Related Crimes", and "Child Sexual Exploitation" for sexual explicit content.
4.3 Evaluation Metrics
Given the critical nature of the task and the importance of accurately identifying and recalling all potentially harmful content, we utilise F1 scores to evaluate our models. We assign equal weight to each label and report the macro-F1 score in both binary and multi-label settings. The F1 scores for individual labels are also considered in the multi-label scenario. This approach helps us gain a deeper insight into the challenges posed by the dataset.
5 Results
The scores for both binary and multi-label scenarios, across all models tested in the fine-tuning and zero-/few-shot settings, are presented in Table 2. In general, fine-tuning leads to clear improvement for all models, which reinforces the importance of our dataset not only to evaluate models, but to build specialised models based based on it. llama3-8b performs best overall, with macro-f1 scores of 85.6 in the binary setting, and 69.8 in the more challenging and fine-grained multi-label setting.
5.1 Fine-tuned Systems
All the fine-tuned models demonstrate high performance with RoBERTa as the least effective, achieving macro-F1 scores of 82.4 in the binary setting and 64.7 in the multi-label setting. The specialised training corpus of tlm appears to enhance its performance, as it consistently surpasses RoBERTa of the same architecture in both settings. Moreover, the fine-tuned version of the larger and more recent llama3-8b model achieves the best overall results in both settings, with macro-F1 scores of 85.6 in the binary setting and 69.8 in the multi-label setting, notably achieved without any hyper-parameter tuning, unlike the other models. Overall, the fine-tuned models tend to struggle the most with the least represented classes, such as Drugs, Selfharm, and Spam. Interestingly, despite comprising 17.3% of the total entries, the models under perform in the Conflictual category, while they exhibit better performance in the less prevalent Sexual Explicit Content class, which accounts for only 9.7% of tweets. This disparity may indicate the models’ difficulties in identifying subtler features within the Conflictual category.
5.2 Zero/Few-shot
Zero-shot.
When tested without any contextual information, the models display varying degrees of effectiveness. Notably, the 70b version of llama3, llama3-70b, outperforms its smaller counterpart and competes with OpenAI’s models, achieving the highest macro F1 score of 76.5 in the binary setting. In the multi-label scenario, the gpt-4o model excels, achieving a macro-F1 score of 64.9. Generally, the zero-shot models do not reach the performance levels of their fine-tuned counterparts, with the notable exception of gpt-4o, which surpasses only RoBERTa in the multi-label setting.
Few-shot.
In the few-shot setting, llama3-8b exhibits performance comparable to its zero-shot execution, illustrating the constraints of smaller models. This limitation is further highlighted by the performance gains observed in the rest of the models, llama3-70b, chat-gpt, and gpt-4o, which show average increases of 7.6 and 3.2 points in macro-F1 for the binary and multi-label settings, respectively. This underscores the effectiveness of in-context learning in larger models. Overall, gpt-4o achieves the best performance, competing with the fine-tuned models and notably outperforming llama3-8b in specific categories, Conflictual, Drugs, and Spam.
5.3 Out of the box
When evaluating the performance of "out-of-the-box" models, we find that they generally fail to achieve high scores (Table 2). Notable exceptions occur in specific categories such as Profanity and Sexual Explicit Content, where the Perspective and llama-g systems excel. The best performing out-of-the-box model, openai-m API, achieves a macro-F1 score of 72% in the binary setting, demonstrating greater robustness in detecting non-sensitive content (F1: 75.9%) compared to its peers. Despite this, its overall performance remains the lowest among the models tested, except for chat-gpt in zero- and few-shot scenarios. It is important to note that these scores may be influenced by the fact that these systems do not utilise the same taxonomy as X-Sensitive, which can impact their performance.
6 Error Analysis
Aiming to understand better the dataset and the challenges that the models face on identifying sensitive content we consider the best performing model, the fine-tuned llama3-8b, and try to understand better its performance.
In the binary setting the model displays a strong performance and achieves high precision and recall values, 85.5%, 85.7% respectively, signifying its ability to effectively identify true positive cases with a relatively low number of false positives (Figure 4, Appendix C.3). In contrast, for the multi-label setting, the model seems to struggle with several categories as seen in Figure 2. Despite the strong performance of llama-8b in classifying Profanity and Sex labels, the model struggles with the Conflictual, Drugs, Self-harm, and Spam categories. As recall increases, precision for these categories drops significantly. This poses a particular challenge for health-related categories like Drugs and Self-harm, where high recall is critical, as missing cases could have serious consequences.
Figure 2: Precision-Recall curve for the fine-tuned llama3-8b in the multi-label setting.
7 Conclusions
In this paper, we presented a complete research approach into sensitive content moderation in social media. Going beyond hate speech, we focus on categories that need to consistently be monitored in social media, let it be to filter to adult users or to remove from the platform, among others. We construct a multi-label dataset using six categories. The results show that LMs fine-tuned on our datasets are generally robust, although there are some categories where they are less precise, and hence these models are probably to be used as a support for human moderators. Nonetheless, the fact that these models perform at a high accuracy represents a useful tool to filter the most relevant messages for each category.
8 Limitations
In this paper, we introduce a valuable new resource expected to benefit a wide range of researchers and industry professionals. However, it is important to acknowledge several limitations. Firstly, the dataset is limited in size, which may restrict the generalisability and robustness of the models trained on it. Additionally, it exclusively contains English-language content due to budget constraints, potentially overlooking the nuances and challenges present in other languages.
The methodology used for aggregating the data in our dataset (Section 3.3) may also be subject to differing opinions. To facilitate transparency and further research, we plan to release all the collected annotations along with the dataset version used in our experiments. Moreover, the dataset was curated based on a specific selection of keywords, which might introduce biases and limit the diversity of the content. Another limitation is that the dataset is derived from only one social media platform, which may not fully represent the variety of sensitive content found across different platforms and contexts.
Finally, while we conduct an in-depth analysis using the results of six different models, there is significant room for improvement in terms of analysis and model development. This includes, investigating the performance of models of different architectures and optimising the prompts used.
9 Ethics Statement
We recognize the significance of the ACL Code of Ethics and are dedicated to adhering to its guidelines in our proposed task. Since our task involves user-generated content, we ensure user privacy by replacing each user mention in the texts with a placeholder, recognising the importance of anonymity, especially taking into account the potential for harm towards people expressing self-harming tendencies.
We also ensure fair treatment of the annotators who labelled the dataset by: 1) compensating them fairly at an average rate of 12\$ per hour, and 2) not sharing or storing personal identification information. As annotator demographics ply an important role in the perception of toxicity, following Prabhakaran et al. (2021), we will release the data, disaggregated by individual annotator labels, while making sure that the demographic information is course enough to prevent deanonymization of the crowdworkers.
Lastly, recognise the sensitive and potentially dangerous nature of the dataset. However, we believe it is crucial to address and combat such behaviours. X-Sensitive will be shared under the CC BY-NC 4.0 Deed (Attribution-NonCommercial 4.0 International) following best practices in sharing social media-based data collections Fiesler and Proferes (2018); Assenmacher et al. (2020).
10 Acknowledgments
A big thank you to Debora Nozza for helpful discussions and feedback on the initial part of this project. Dimosthenis Antypas and Jose Camacho-Collados are supported by a UKRI Future Leaders Fellowship.
Appendix A Data Collection
Table 3 displays the sources used for finding keywords used to collect tweets.
Category Sources
Selfharm Chancellor et al. (2016b)
Chancellor et al. (2016d)
Chancellor et al. (2016a)
Chancellor et al. (2016c)
Ngrams from the Reddit training corpus
Conflictual, profane, sexually explicit IDEA-NTHU-Taiwan
LDNOOBW
Drugs talktofrank
Guns Wikipedia
Spam Founta et al. (2018)
#f4f, #l4l, #follow4follow, #like4like
Table 3: Sources for seed words used for expanding our keyword lists.
Table 4 displays the total number of keywords and see words used in our data collection.
Category #seed-words #keywords
Conflictual,
profane,
sexually explicit
203 1322
spam 76 767
drugs 118 519
selfharm 56 734
Table 4: Overall summary of the keyword list.
Appendix B Annotation
B.1 Guidelines
Figure 3 displays the guidelines provide to each coder for the annotation task.
Figure 3: Guidelines provided to annotators.
B.2 Annotator Agreement
Table 5 displays the inter annotator agreement based on Krippendorff’s alpha between different gender groups of coders. Similarly, Table 6 displays agreement scores between different codes age groups.
Gender Multi Bin
Man 0.49 0.57
Woman 0.49 0.55
Non-binary 0.47 0.82
Table 5: Krippendorff’s alpha withing each gender group of coders.
Age Multi Bin
18-25 0.51 0.57
26-39 0.51 0.59
40-65 0.47 0.52
over 65 0.47 0.54
Table 6: Krippendorff’s alpha withing each age group of coders.
Tables 7 and 8 show the percentage of tweets labelled as each class by age and gender groups of coders, respectively.
Age conflictual profanity sex drugs selfharm spam AVG Coders
18-25 15 28 9 3 3 4 10 77
26-39 13 28 9 3 3 4 10 269
40-65 15 25 8 3 2 3 9 166
over 65 10 27 5 3 1 2 8 10
Table 7: Percentage of tweets labelled as each class for each age bracket of coders.
Gender conflictual profanity sex drugs selfharm spam AVG Coders
Woman 14 27 8 3 3 3 10 268
Non-Binary 20 28 10 2 3 3 11 12
Man 14 26 9 3 3 4 10 241
Table 8: Percentage of tweets labelled as each class for each gender of coders.
Appendix C Models
C.1 Resources
In total we estimate 112 hours used for the training of RoBERTa, tlm and llama3-8b models using a NVIDIA GeForce RTX 4090 GPU and 90 hours for inferences with the llama3-8b and llama3-70b models using an NVIDIA Quadro RTX 8000 GPU. Table 9 provides details for the models used in our experiments.
Model Parameters
RoBERTa 355M
tlm 355M
llama3-8b 8B
llama3-70b 70B
chat-gpt 175B (approximate)
Table 9: Number of Parameters in different language models used.
C.2 Taxonomies
Taxonomy used by Google’s perspective API:
  1. 1.
    TOXICITY
  2. 2.
    SEVERE_TOXICITY
  3. 3.
    IDENTITY_ATTACK
  4. 4.
    INSULT
  5. 5.
    PROFANITY
  6. 6.
    SEXUALLY_EXPLICIT
  7. 7.
    THREAT
  8. 8.
    FLIRTATION
  9. 9.
    ATTACK_ON_AUTHOR
  10. 10.
    ATTACK_ON_COMMENTER
  11. 11.
    INCOHERENT
  12. 12.
    INFLAMMATORY
  13. 13.
    LIKELY_TO_REJECT
  14. 14.
    OBSCENE
  15. 15.
    SPAM
  16. 16.
    UNSUBSTANTIAL
In our experiments we utilise the following class mapping to the X-Sensitive taxonomy: "TOXICITY": Conflictual, "PROFANITY": Profanity, "SEXUALLY_EXPLICIT": Sexual Explicit Content, "SPAM": Spam.
MLCommons taxonomy used in Meta-LLama-guard:
  • 1:
    Violent Crimes
  • 2:
    Non-Violent Crimes
  • 3:
    Sex-Related Crimes
  • 4:
    Child Sexual Exploitation
  • 5:
    Specialized Advice
  • 6:
    Privacy
  • 7:
    Intellectual Property
  • 8:
    Indiscriminate Weapons
  • 9:
    Hate
  • 10:
    Suicide & Self-Harm
  • 11:
    Sexual Content
Openai’s moderation API taxonomy:
  • 1:
    harassment
  • 2:
    harassment_threatening
  • 3:
    hate
  • 4:
    hate_threatening
  • 5:
    self_harm
  • 6:
    self_harm_instructions
  • 7:
    self_harm_intent
  • 8:
    sexual
  • 9:
    sexual_minors
  • 10:
    violence
  • 11:
    violence_graphic
  • 12:
    self-harm
  • 13:
    sexual/minors
  • 14:
    hate/threatening
  • 15:
    violence/graphic
  • 16:
    self-harm/intent
  • 17:
    self-harm/instructions
  • 18:
    harassment/threatening
In our experiments we utilise the following class mapping to the X-Sensitive taxonomy:
’harassment’: ’Conflictual’,
’harassment_threatening’: ’Conflictual’,
’hate’: ’Conflictual’,
’hate_threatening’: ’Conflictual’,
’self_harm’: ’Self Harm’,
’self_harm_instructions’: ’Self Harm’,
’self_harm_intent’: ’Self Harm’,
’sexual’: ’Sexual Explicit Content’,
’sexual_minors’: ’Sexual Explicit Content’,
’self-harm’: ’Self Harm’,
’sexual/minors’: ’Sexual Explicit Content’,
’hate/threatening’: ’Conflictual’,
’self-harm/intent’: ’Self Harm’,
’self-harm/instructions’: ’Self Harm’,
’harassment/threatening’: ’Conflictual’
C.3 Extended Results
Figure 4 displays the Precision-Recall curve for the fine-tuned llama3-8b model in the binary setting.
Figure 4: Precision-Recall curve for the fine-tuned llama3-8b model in the binary setting.